A group of students got the following scores in an achievement test: 9, 12 15, 18, 21 and 24. Consider samples of size 3 can be drawn from this population.


B. Construct the sampling distribution of the resulting sample means



1
Expert's answer
2022-04-20T10:55:52-0400

The number of possible samples of size n=3 from a population of size N=6 is,

"=C_6^3= \binom{6}{3}=20"


The sample means are derived from the formula, "\bar x_i={\sum x_i\over3}"



The sampling distribution is as follows:

"\begin{matrix}{\bar x_i}&{12}&{13}&{14}&{15}&{16}&{17}&{18}&{19}&{20}&{21}\\{p(\bar x_i)}&{\frac{1}{{20}}}&{\frac{1}{{20}}}&{\frac{1}{{10}}}&{\frac{3}{20}}&{\frac{3}{{20}}}&{\frac{3}{20}}&{\frac{3}{{20}}}&{\frac{1}{10}}&{\frac{1}{{20}}}&{\frac{1}{{20}}}\end{matrix}"